# Free-form language-based robotic reasoning and grasping

### v1

@@ Abstract
Performing robotic grasping from a cluttered bin based on human instructions is a challenging task, as it requires understanding both the nuances of free-form language and the spatial relationships between objects. Vision-Language Models (VLMs) trained on web-scale data, such as GPT-4o, have demonstrated remarkable reasoning capabilities across both text and images. But can they truly be used for this task in a zero-shot setting? And what are their limitations? In this paper, we explore these research questions via the free-form language-based robotic grasping task, and propose a novel method, FreeGrasp, leveraging the pre-trained VLMs’ world knowledge to reason about human instructions and object spatial arrangements. Our method detects all objects as keypoints and uses these keypoints to annotate marks on images, aiming to facilitate GPT-4o’s zero-shot spatial reasoning. This allows our method to determine whether a requested object is directly graspable or if other objects must be grasped and removed first. Since no existing dataset is specifically designed for this task, we introduce a synthetic dataset FreeGraspData by extending the MetaGraspNetV2 dataset with human-annotated instructions and ground-truth grasping sequences. We conduct extensive analyses with both FreeGraspData and real-world validation with a gripper-equipped robotic arm, demonstrating state-of-the-art performance in grasp reasoning and execution. Project website: https://tev-fbk.github.io/FreeGrasp/.
@@ I Introduction
Vision-Language Models (VLMs) encode vast semantic knowledge about the world, which is essential for interpreting nuanced human free-form language instructions [1]. This valuable source of information can enable robots to operate effectively in diverse, unseen, and unstructured environments [2, 3]. Interpreting human instructions is only one part of the challenge; grounding the retrieved information in the real world is equally critical. This process requires spatial reasoning [4, 5], i.e. understanding and acting upon spatial relationships between objects [6, 7]. By achieving these capabilities together, we can tackle novel problems and enable embodied agents to perform object manipulation and rearrangement in less-controlled environments [8, 9, 10, 11].
Recently, several studies have explored the use of pre-trained large-scale models for robotic control [12, 13, 14]. Building on this direction, we investigate how to leverage VLMs’ knowledge to enable a robot to interpret human commands. However, unlike previous works, our focus is specifically on evaluating the robustness of VLMs to free-form language instructions combined with spatial reasoning capabilities for robotic grasping. Our setup, illustrated in Fig. 1, consists of a robotic arm with a two-finger gripper, an exocentric RGB-D camera, a red bin containing various objects in clutter, and an empty blue bin where the grasped objects are put. The objects in the red bin are arranged haphazardly, and multiple instances of the same object may be present, e.g. Rubik’s cubes in this case. Our robotic system is commanded by users through instructions, e.g. grasp the Rubik’s cube on the top right, with the goal of identifying which objects must be removed first in order to grasp the user-specified target. The system should minimize the number of grasping actions to complete the task efficiently.
@@ II Related works
VLM-driven robotic grasping. Multimodal learning has significantly enhanced robots’ ability to interpret language instructions and perform grasp reasoning tasks. Research has focused on task-oriented grasping and spatial affordance prediction through VLMs. For example, RoboPoint [17] integrates VLMs to predict spatial affordances, improving object interaction understanding. SpatialVLA [18] enhances Vision-Language-Action (VLA) models with spatial representations, improving task planning and execution. GLOVER [19] advances open-vocabulary reasoning for task-oriented grasping, enabling robots to perform diverse tasks without task-specific training. To improve spatial reasoning, SpatialCoT [20] combines coordinate alignment with chain-of-thought reasoning for improved embodied task planning. SpatialPIN [21] refines this by integrating 3D priors and advanced prompting techniques, enhancing object recognition, scene understanding, and navigation. Authors in [22, 23, 14, 24] focus on enabling robots to follow natural language instructions for manipulation tasks. ThinkGrasp [14], somewhat related to our work, combines visual recognition and language-based reasoning to improve part grasping in cluttered environments. Unlike ThinkGrasp, our method, FreeGrasp, can handle free-form language instructions perform spatial reasoning more effectively. Moreover, FreeGrasp can interpret complex language descriptions to resolve object ambiguities, ensuring accurate identification of target objects even in cluttered or ambiguous environments.
@@ III-A Overview
We address the robotic reasoning and grasping task, and propose an effective and modular pipeline by leveraging the world-knowledge of pre-trained VLMs without additional training. Our setup consists of a bin containing objects organized haphazardly, an exocentric RGB-D camera with top-down view, a robotic manipulator, and a user who provides a free-form language instruction to request the robot to grasp a target object, e.g. “grasp the red toy car at the bottom.” Fig. 2 shows the diagram of FreeGrasp.
An episode starts upon receiving the user’s instruction. At each step, the pipeline starts with the proposed vision-language grasp reasoning module. With the RGB observation, the module first localizes all the objects within the container, forming a holistic understanding of the scene [32]. To facilitate VLMs in visual spatial reasoning, we augment the visual prompt by annotating identity (ID) marks for each localized object on the input image. Then, we feed the mark-based visual prompt to GPT-4o [1] to reason about object spatial relationships. GPT-4o determines whether the target object is free of obstruction, thus directly graspable, or is obstructed by other objects that must be removed first. GPT-4o’s output consists of the ID and class name of the next object to be grasped, along with information indicating whether it is the target object. With the class name and ID, the object segmentation model further segments the output object [33]. The grasp estimation module determines the most suitable grasp pose for picking the segmented object [15]. Lastly, the robotic arm plans and executes its trajectory to grasp the object and place it in a predefined location, e.g., another bin. After each step, the positions of objects may change. The pipeline thus repeats with a new RGB-D acquisition. The episode for the robotic reasoning and grasping task terminates until the user-specified target object is being grasped, or certain criteria is met.
@@ III-B Vision-language grasp reasoning
Given the RGB image capturing the scene and the user’s free-form language instruction, this module leverages state-of-the-art VLMs to visually ground objects and reason on their relationship to determine the object to grasp, without additional model training.
@@
Figure 3: Object localization performance with different VLM-based method on MetaGrasNetv2 [16].
@@
Object segmentation. We segment the object instance identified in the image. For instance segmentation, simply using the 2D coordinate given by the object ID as visual prompt to class-agnostic segmentation models, e.g. SAM [39], results in inaccurate instance mask with little semantic control. We thus first perform semantic segmentation using LangSAM [33] (Fig. 2) with the class name. As there might be multiple instances with the same class, we then use the object ID to filter the instance mask of interest.
@@ III-C Grasp estimation
IV Free-form language grasping dataset
We introduce the free-from language grasping dataset (FreeGraspData), a novel dataset built upon MetaGraspNetv2 [16] to evaluate the robotic grasping task with free-form language instructions. MetaGraspNetv2 is a large-scale simulated dataset featuring challenging aspects of robot vision in the bin-picking setting, including multi-view RGB-D images and metadata, e.g. object categories, amodal segmentation masks, and occlusion graphs indicating occlusion relationships between objects from each viewpoint. To build FreeGraspData, we selected scenes containing at least four objects to ensure sufficient scene clutter. FreeGraspData extends MetaGraspNetV2 in three aspects: i) we derive the ground-truth grasp sequence until reaching the target object from the occlusion graphs, ii) we categorize the task difficulty based on the obstruction level and instance ambiguity, and iii) we provide free-form language instructions, collected from human annotators.
IV-A Ground-truth grasp sequence
We obtain the ground-truth grasp sequence based on the object occlusion graphs provided in MetaGraspNetV2. As the visual occlusion does not necessarily indicate obstruction, we thus first prune the edges in the provided occlusion graph that are less likely to form obstruction. Following the heuristic that less occlusion indicates less chance of obstruction, we remove the edges where the percentage of the occlusion area of the occluded object is below $1\%$. From the node representing the target object, we can then traverse the pruned graph to locate the leaf node, that is the ground-truth object to grasp first. The sequence from the leaf node to the target node forms the correct sequence for the robotic grasping task.
IV-B Grasp difficulty categorization
IV-C Free-form language user instructions
For each of the six difficulty categories, we randomly select 50 objects, resulting in 300 robotic grasping scenarios. For each scenario, we provide multiple users with a top-down image of the bin and a visual indicator highlighting the target object. No additional context or information about the object is provided. We instruct the user to provide an unambiguous natural language description of the indicated object with their best effort. In total, ten users are involved in the data collection procedure, with a wide age span. We randomly select three user instructions for each scenario, yielding a total of 900 evaluation scenarios. This results in diverse language instructions, as shown in Fig. 4.
@@ IV-C Free-form language user instructions
V Experiments
We conduct experiments using both our synthetic evaluation dataset, FreeGraspData, and a gripper-equipped robotic arm in the real world. We compare FreeGrasp against the state-of-the-art method, ThinkGrasp [14], which also uses GPT-4o for reasoning. To comprehensively evaluate robotic grasping, we introduce novel metrics that assess performance based on both intermediate steps and the final robotic grasp. Using FreeGraspData, we also perform extensive ablation studies on the key components of FreeGrasp to validate our design choices. Lastly, through real-world experiments, we demonstrate the effectiveness of FreeGrasp in handling practical challenges such as imperfect depth measurements and robotic grasp execution. More robotic demonstrations can be found in the Supplementary Material.
V-A Experiments on FreeGraspData
Performance metrics. Since FreeGraspData is a static dataset, we mainly evaluate grasp reasoning and object segmentation. We do not assess grasp estimation with FreeGraspData, as this aspect is best evaluated through robotic execution (Sec. V-B). For grasp reasoning, we report the Reasoning Success Rate (RSR), i.e. the proportion of successful episodes in which the predicted object ID matches one of the ground-truth IDs. For object segmentation, we report the Segmentation Success Rate (SSR), i.e. the proportion of successful episodes in which the output mask achieves an Intersection over Union of at least 0.5 with the ground-truth mask. Since ThinkGrasp does not produce object IDs, we can only report its SSR. For FreeGrasp, we report both RSR and SSR. Moreover, we evaluate a variant of FreeGrasp to assess the impact of object localization on reasoning, as measured by RSR. This variant uses ground-truth (GT) localization instead of Molmo.
[…]
V-B Experiments in the real world
Experimental setup. We evaluate FreeGrasp and ThinkGrasp [14] in a real-world scenario with a UR5e robotic arm equipped with an OnRobot RG2 parallel gripper. The scene setup is similar as in prior work ThinkGrasp [14] for fair comparison. Specifically, the robotic arm is in a table-top setting. The scene is captured by a RealSense D415 RGB-D camera with a top-down view featuring the cluttered red bin, at a resolution of 1280$\times$720. We implement the motion planning and control via the MoveIt motion planning framework with ROS. Adhering to the definition of six task difficulties in Sec. IV, we compose ten evaluation scenarios per difficulty level (Fig. 6), each featuring a unique arrangement of objects (e.g. screwdrivers, stationery items, boxes, toys, etc) with a single free-form language instruction.
To study the impact of localized failures, we further devise three operational settings by defining the stopping criteria based on localized failures: i) Setting (S, P, M) is the most stringent setting, where an evaluation episode will be terminated if any of the localized failure occurs throughout the robotic reasoning and grasping task; ii) Setting (P, M) relaxes the segmentation failure, where an evaluation episode will be terminated only if a pose failure or motion failure occurs; and iii) Setting (P) relaxes the motion failure, where an evaluation episode will only be terminated if a pose failure occurs. If motion failure under Setting (P) occurs, we manually move the object to the bin. If no failure occurs, the episode continues until the target object is being grasped and placed.
Performance metrics. We use three metrics, inspired by navigation tasks [42], to measure the effectiveness and efficiency of ThinkGrasp and FreeGrasp. We define the success of an episode when the correct target object is grasped and placed to its pre-defined destination. The success rate (SR) reports the ratio of successful episodes among all evaluation episodes, i.e. $SR=\frac{N_{s}}{N}$, where $N$ is the number of total evaluation episodes and $N_{s}$ is the number of successful episodes. The path efficiency (SE) reports the normalized inverse step counts, i.e. $PE=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\frac{l_{i}}{p_{i}}$, where $l_{i}$ is the number of minimum steps to solve the episode as defined by an oracle (i.e. the operator), and $p_{i}$ is the actual number of steps the robot actually had to take to solve the episode. Finally, the Success Weighted (normalized inverse) path length (SPL), calculated as $SPL=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{p_{i}}$, where $S_{i}$ is $1$ if episode $i$ is successful, 00 otherwise.
@@ VI Conclusions
We introduced FreeGrasp, a novel approach that leverages pre-trained VLMs for robotic grasping by interpreting free-form instructions and reasoning about spatial relationships. Our investigation showed that while VLMs, e.g. GPT-4o, are known for strong general reasoning capabilities, they struggle with visual spatial reasoning, highlighting a important gap for both visual grounding and spatial awareness. FreeGrasp was designed to mitigate such gap with mark-based visual prompting and contextualized reasoning, outperforming the state-of-the-art, ThinkGrasp, in both synthetic and real-world robotic validations. Overall, FreeGrasp demonstrates the potential of combining VLMs with modular reasoning to tackle robotic grasping challenges in real-world scenarios.
Limitations and future works. A key limitation we observed of FreeGrasp is GPT-4o’s limited visual-spatial capability, particularly in understanding object occlusion While we tested specialized spatial VLMs like SpaceLLaVA [6], they underperformed compared to GPT-4o. Another limitation is the lack of a mechanism to adapt the initial human instructions as scenes change with object removal. For example, prompts like the duck on the right of the screwdriver become invalid if the screwdriver is removed. Future work could explore memory mechanisms to track scene changes or adaptive instruction updates using VLMs to dynamically adjust references as objects are removed.

### v2

@@ Abstract
Performing robotic grasping from a cluttered bin based on human instructions is a challenging task, as it requires understanding both the nuances of free-form language and the spatial relationships between objects. Vision-Language Models (VLMs) trained on web-scale data, such as GPT-4o, have demonstrated remarkable reasoning capabilities across both text and images. But can they truly be used for this task in a zero-shot setting? And what are their limitations? In this paper, we explore these research questions via the free-form language-based robotic grasping task and propose a novel method, FreeGrasp, leveraging the pre-trained VLMs’ world knowledge to reason about human instructions and object spatial arrangements. Our method detects all objects as keypoints and uses these keypoints to annotate marks on images, aiming to facilitate GPT-4o’s zero-shot spatial reasoning. This allows our method to determine whether a requested object is directly graspable or if other objects must be grasped and removed first. Since no existing dataset is specifically designed for this task, we introduce a synthetic dataset FreeGraspData by extending the MetaGraspNetV2 dataset with human-annotated instructions and ground-truth grasping sequences. We conduct extensive analyses with both FreeGraspData and real-world validation with a gripper-equipped robotic arm, demonstrating state-of-the-art performance in grasp reasoning and execution. Project website: https://tev-fbk.github.io/FreeGrasp/.
@@ I Introduction
Vision-Language Models (VLMs) encode vast semantic knowledge about the world, which is essential for interpreting nuanced human free-form language instructions [1]. This valuable source of information can enable robots to operate effectively in diverse, unseen, and unstructured environments [2, 3]. Interpreting human instructions is only one part of the challenge; grounding the retrieved information in the real world is equally critical. This process requires spatial reasoning [4, 5], i.e. understanding and acting upon spatial relationships between objects [6, 7]. By achieving these capabilities together, we can tackle novel problems and enable embodied agents to perform object manipulation and rearrangement in less-controlled environments [8].
Recently, several studies have explored the use of pre-trained large-scale models for robotic control [9, 10, 11]. Building on this direction, we investigate how to leverage VLMs’ knowledge to enable a robot to interpret human commands. However, unlike previous works, our focus is specifically on evaluating the robustness of VLMs to free-form language instructions combined with spatial reasoning capabilities for robotic grasping. Our setup, illustrated in Fig. 1, consists of a robotic arm with a two-finger gripper, an exocentric RGB-D camera, a red bin containing various objects in clutter, and an empty blue bin where the grasped objects are put. The objects in the red bin are arranged haphazardly, and multiple instances of the same object may be present, e.g., Rubik’s cubes in this case. Our robotic system is commanded by users through instructions, e.g. grasp the Rubik’s cube on the top right, with the goal of identifying which objects must be removed first in order to grasp the user-specified target. The system should minimize the number of grasping actions to complete the task efficiently.
@@ II Related works
VLM-driven robotic grasping. Multimodal learning has significantly enhanced robots’ ability to interpret language instructions and perform grasp reasoning tasks. Research has focused on task-oriented grasping and spatial affordance prediction through VLMs. For example, RoboPoint [14] integrates VLMs to predict spatial affordances, improving object interaction understanding. SpatialVLA [15] enhances Vision-Language-Action (VLA) models with spatial representations, improving task planning and execution. GLOVER [16] advances open-vocabulary reasoning for task-oriented grasping, enabling robots to perform diverse tasks without task-specific training. To improve spatial reasoning, SpatialCoT [17] combines coordinate alignment with chain-of-thought reasoning for improved embodied task planning. SpatialPIN [18] refines this by integrating 3D priors and advanced prompting techniques, enhancing object recognition, scene understanding, and navigation. ThinkGrasp [11], somewhat related to our work, combines visual recognition and language-based reasoning to improve part grasping in cluttered environments. Unlike ThinkGrasp, our method, FreeGrasp, can handle free-form language instructions perform spatial reasoning more effectively. Moreover, FreeGrasp can interpret complex language descriptions to resolve object ambiguities, ensuring accurate identification of target objects even in cluttered or ambiguous environments.
@@ III-A Overview
We address the robotic reasoning and grasping task and propose an effective and modular pipeline by leveraging the world knowledge of pre-trained VLMs without additional training. Our setup consists of a bin containing objects organized haphazardly, an exocentric RGB-D camera with top-down view, a robotic manipulator, and a user who provides a free-form language instruction to request the robot to grasp a target object, e.g. “grasp the red toy car at the bottom.” Fig. 2 shows the diagram of FreeGrasp.
An episode starts upon receiving the user’s instruction. At each step, the pipeline starts with the proposed vision-language grasp reasoning module. With the RGB observation, the module first localizes all the objects within the container, forming a holistic understanding of the scene [26]. To facilitate VLMs in visual spatial reasoning, we augment the visual prompt by annotating identity (ID) marks for each localized object on the input image. Then, we feed the mark-based visual prompt to GPT-4o [1] to reason about object spatial relationships. GPT-4o determines whether the target object is free of obstruction, thus directly graspable, or is obstructed by other objects that must be removed first. GPT-4o’s output consists of the ID and class name of the next object to be grasped, along with information indicating whether it is the target object. With the class name and ID, the object segmentation model further segments the output object [27]. The grasp estimation module determines the most suitable grasp pose for picking the segmented object [12]. Lastly, the robotic arm plans and executes its trajectory to grasp the object and place it in a predefined location, e.g., another bin. After each step, the positions of objects may change. The pipeline thus repeats with a new RGB-D acquisition. The episode for the robotic reasoning and grasping task terminates until the user-specified target object is being grasped or certain criteria is met.
@@ III-B Vision-language grasp reasoning
Given the RGB image capturing the scene and the user’s free-form language instruction, this module leverages state-of-the-art VLMs to visually ground objects and reason about their relationship to determine the object to grasp, without additional model training.
@@
Figure 3: Object localization performance with different VLM-based method on MetaGrasNetv2.
[…]
Figure 4: Occlusion graphs in representative scenarios. (a) Real-world scene and its corresponding occlusion graph. (b) Synthetic scene from FreeGraspData and its occlusion graph. Objects are colored by difficulty, and directed edges indicate occlusion dependencies. Example task instructions and step-by-step action plans are shown beneath each graph.
@@
Object segmentation. We segment the object instance identified in the image. For instance segmentation, simply using the 2D coordinate given by the object ID as visual prompt to class-agnostic segmentation models, e.g. SAM [33], results in inaccurate instance masks with little semantic control. We thus first perform semantic segmentation using LangSAM [27] (Fig. 2) with the class name. As there might be multiple instances with the same class, we then use the object ID to filter the instance mask of interest.
@@ III-C Grasp estimation
IV Free-form language grasping dataset
We introduce the free-from language grasping dataset (FreeGraspData), a novel dataset built upon MetaGraspNetv2 [13] to evaluate the robotic grasping task with free-form language instructions. MetaGraspNetv2 is a large-scale simulated dataset featuring challenging aspects of robot vision in the bin-picking setting, including multi-view RGB-D images and metadata, e.g. object categories, amodal segmentation masks, and occlusion graphs indicating occlusion relationships between objects from each viewpoint. To build FreeGraspData, we selected scenes containing at least four objects to ensure sufficient scene clutter. FreeGraspData extends MetaGraspNetV2 in three aspects: i) we derive the ground-truth grasp sequence until reaching the target object from the occlusion graphs, ii) we categorize the task difficulty based on the obstruction level and instance ambiguity, and iii) we provide free-form language instructions collected from human annotators. Examples of occlusion graphs for real-world and FreeGraspData scenes are illustrated in Fig. 4.
IV-A Ground-truth grasp sequence
[…]
IV-B Grasp difficulty categorization
IV-C Free-form language user instructions
For each of the six difficulty categories, we randomly select 50 objects, resulting in 300 robotic grasping scenarios. For each scenario, we provide multiple users with a top-down image of the bin and a visual indicator highlighting the target object. We randomly select three user instructions for each scenario, yielding a total of 900 evaluation scenarios. This results in diverse language instructions, as shown in Fig. 5.
@@ IV-C Free-form language user instructions
V Experiments
We conduct experiments using both our synthetic evaluation dataset, FreeGraspData, and a gripper-equipped robotic arm in the real world. We compare FreeGrasp against the state-of-the-art method, ThinkGrasp [11], which also uses GPT-4o for reasoning. To comprehensively evaluate robotic grasping, we introduce novel metrics that assess performance based on both intermediate steps and the final robotic grasp. Using FreeGraspData, we also perform extensive ablation studies on the key components of FreeGrasp to validate our design choices. Lastly, through real-world experiments, we demonstrate the effectiveness of FreeGrasp in handling practical challenges such as imperfect depth measurements and robotic grasp execution.
V-A Experiments on FreeGraspData
Performance metrics. Since FreeGraspData is a static dataset, we mainly evaluate grasp reasoning and object segmentation. We do not assess grasp estimation with FreeGraspData, as this aspect is best evaluated through robotic execution (Sec. V-B). For grasp reasoning, we report the Reasoning Success Rate (RSR), i.e. the proportion of successful episodes in which the predicted object ID matches one of the ground-truth IDs. For object segmentation, we report the Segmentation Success Rate (SSR), i.e. the proportion of successful episodes in which the output mask achieves an Intersection over Union of at least 0.5 with the ground-truth mask. Since ThinkGrasp does not produce object IDs, we can only report its SSR. Moreover, we evaluate a variant of FreeGrasp to assess the impact of object localization on reasoning, as measured by RSR. This variant uses ground-truth (GT) localization instead of Molmo.
V-B Experiments in the real world
Experimental setup. We evaluate FreeGrasp and ThinkGrasp [11] in a real-world scenario with a UR5e robotic arm equipped with an OnRobot RG2 parallel gripper. The scene setup is similar as in prior work ThinkGrasp [11] for fair comparison. The scene is captured by a RealSense D415 RGB-D camera with a top-down view featuring the cluttered red bin, at a resolution of 1280$\times$720. We implement the motion planning and control via the MoveIt motion planning framework with ROS. Adhering to the definition of six task difficulties in Sec. IV, we compose ten evaluation scenarios per difficulty level (Fig. 7), each featuring a unique arrangement of objects
To study the impact of localized failures, we further devise three operational settings by defining the stopping criteria based on localized failures: i) Setting (S, P, M) is the most stringent setting, where an evaluation episode will be terminated if any of the localized failures occur throughout the robotic reasoning and grasping task; ii) Setting (P, M) relaxes the segmentation failure, where an evaluation episode will be terminated only if a pose failure or motion failure occurs; and iii) Setting (P) relaxes the motion failure, where an evaluation episode will only be terminated if a pose failure occurs.
Performance metrics. We use three metrics, inspired by navigation tasks [36], to measure the effectiveness and efficiency of ThinkGrasp and FreeGrasp. We define the success of an episode when the correct target object is grasped and placed to its pre-defined destination. The success rate (SR) reports the ratio of successful episodes among all evaluation episodes, i.e. $SR=\frac{N_{s}}{N}$, where $N$ is the number of total evaluation episodes and $N_{s}$ is the number of successful episodes. The path efficiency (SE) reports the normalized inverse step counts, i.e. $PE=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\frac{l_{i}}{p_{i}}$, where $l_{i}$ is the number of minimum steps to solve the episode as defined by an oracle (i.e. the operator), and $p_{i}$ is the actual number of steps the robot actually had to take to solve the episode. Finally, the Success Weighted (normalized inverse) path length (SPL), calculated as $SPL=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{p_{i}}$, where $S_{i}$ is 111 if episode $i$ is successful, 0 otherwise.
@@ VI Conclusions
We introduced FreeGrasp, a novel approach that leverages pre-trained VLMs for robotic grasping by interpreting free-form instructions and reasoning about spatial relationships. Our investigation showed that while VLMs, e.g. GPT-4o, are known for strong general reasoning capabilities, they struggle with visual spatial reasoning, highlighting an important gap for both visual grounding and spatial awareness. FreeGrasp was designed to mitigate such gap with mark-based visual prompting and contextualized reasoning, outperforming the state-of-the-art, ThinkGrasp, in both synthetic and real-world robotic validations. Overall, FreeGrasp demonstrates the potential of combining VLMs with modular reasoning to tackle robotic grasping challenges in real-world scenarios.
Limitations and future works. A key limitation we observed of FreeGrasp is GPT-4o’s limited visual-spatial capability, particularly in understanding object occlusion. While we tested specialized spatial VLMs like SpaceLLaVA [6], they underperformed compared to GPT-4o. Another limitation is the lack of a mechanism to adapt the initial human instructions as scenes change with object removal. Future work could explore memory mechanisms to track scene changes or adaptive instruction updates using VLMs to dynamically adjust references as objects are removed.